# Counterexamples to maximal regularity for operators in divergence form

### v1

@@ 1. Introduction
(CP) {$u^{\prime}-\mathcal{A}u&=f,\ \ \text{on}\ (0,1),\\ u(0)&=0.\end{aligned}\right.$
@@
(P) {$u^{\prime}-{\text{div}}(B\nabla u)&=f,\quad\text{on }(0,1),\\ u(0)&=0,\end{aligned}\right.$
@@ 2.1. Known results
[…]
@@ 2.2. Failure of $L^{r}(L^{s})$-integrability for variational solutions
The following counterexample is based on a construction due to the second-named author [19] and is the basis for our subsequent counterexamples to Problems 1 and 2.
@@
Pick $\frac{2}{r}+\frac{d}{s}<\mu<\frac{d}{2}$.
Step 1: general setup. We repeat part of the construction in [19]. There exist a Lipschitz function $w:{\mathbb{R}}^{d}\to{\mathbb{C}}$ and an elliptic matrix $a:{\mathbb{R}}^{d}\to{\mathbb{C}}^{d\times d}$ with the subsequent properties. The functions $a$ and $w$ are uniformly bounded, smooth away from zero, Lipschitz in zero, and for all multiindices $\alpha$, $|x|\geq 1$,
(2.2) $|\partial^{\alpha}w(x)|$ $\leq C_{\alpha}|x|^{-|\alpha|-\mu},$
(2.3) $|\partial^{\alpha}a(x)|$ $\leq C_{\alpha}|x|^{-|\alpha|}.$
These estimates follow from the definitions of $w$ and $a$ given in [19, Sec. 3]. Similar as in [19, Eq. (5), Thm 2.2, and Cor 2.3] define
$B(t,x)$ $\coloneqq a(x/(1-t)^{1/2})$
$\zeta(t,x)$ $\coloneqq(1-t)^{-\mu/2}e^{-\frac{i}{2}\log(1-t)}w(x/(1-t)^{1/2}).$
These were constructed so that $\zeta\in L^{2}_{\rm loc}([0,1)\times{\mathbb{R}}^{d})$, $\nabla\zeta\in L^{2}_{\rm loc}([0,1)\times{\mathbb{R}}^{d})$ and the equation
(2.4) $\partial_{t}\zeta(t,x)={\text{div}}(B(t,x)\nabla\zeta(t,x))$
is satisfied pointwise for $t\in(0,1)$ and $x\in{\mathbb{R}}^{d}\setminus\{0\}$. The latter means that
$(\zeta,\partial_{t}\phi)_{L^{2}((0,1)\times{\mathbb{R}}^{d})}=(B\nabla\zeta,\nabla\phi)_{L^{2}((0,1)\times{\mathbb{R}}^{d})}$
holds for all $\phi\in C^{\infty}_{c}((0,1)\times{\mathbb{R}}^{d})$. It is clear from the constructions that the weak derivative $\partial_{t}\zeta$ exists on $(0,T)$ as an $L^{2}(B_{R})$-valued function, for any $T<1$ and $R>0$. Now let
@@
$\partial_{t}u(t,x)-{\text{div}}(B(t,x)\nabla u(t,x))&=f(t,x),\\ u(t,0)&=0.\end{aligned}\right.$
@@
By choice of $\mu$ we have $-\frac{\mu+1}{2}+\frac{d}{4}>-\frac{1}{2}$, therefore also $\nabla u\in L^{2}(0,1;L^{2}({\mathbb{R}}^{d}))$.
@@
(2.6) $\begin{split}\|\partial_{m}\zeta(t,\cdot)\|_{L^{2}(B_{2}\setminus B_{1})}&=(1-t)^{-\frac{\mu}{2}-\frac{1}{2}}\|\partial_{m}w(\cdot/(1-t)^{1/2}\|_{L^{2}(B_{2}\setminus B_{1})}\\ &\leq C_{1}\||\cdot|^{-1-\mu}\|_{L^{2}(B_{2}\setminus B_{1})}.\end{split}$
@@
By choice of $\mu$, $r$ and $s$ we have $-\frac{\mu}{2}+\frac{d}{2s}<-\frac{1}{r}$, therefore $u\not\in L^{r}(0,1;L^{s}({\mathbb{R}}^{d}))$. ∎
@@ 2.3. Negative result concerning problem 1
(2.7) $\partial_{t}u-{\text{div}}(B\nabla u)&=f,\\ u(0)&=0,\end{aligned}\right.$
@@
(2.8) $\partial_{t}\widetilde{v}-{\text{div}}(A\nabla\widetilde{v})&=\widetilde{g},\\ \widetilde{v}(-1)&=0,\end{aligned}\right.$
@@
where the constant $C>0$ does not depend on $\widetilde{v}$ and $\widetilde{g}$. By translation the question on $(-1,0)$ is equivalent to that on $(0,1)$. Both intervals are related by the transformation $v\to-v$. We write for instance $\widetilde{u}(t)=u(-t)$ to translate $u$ to a function on $(-1,0)$ and vice versa. Specialize $A=(\widetilde{B})^{*}$ in (2.8). Since $u(0)=0=v(1)$ we can use integration by parts to obtain
@@
[…]
@@ 2.4. Negative result concerning problem 2
The strategy is as for the case $p>2$ in Theorem 2.4. Let $\nu\in(1/2,1]$ and let $0<\theta<\frac{1}{2\nu}$. Then define parameters $r$ and $s$ through $\frac{1}{r}=\frac{1}{2}-\nu\theta$ and $\frac{1}{s}=\frac{1}{2}-\frac{1-\theta}{d}$. They satisfy the condition in Theorem 2.4, consequently there is $B:[0,1]\times{\mathbb{R}}^{d}\to{\mathbb{C}}^{d\times d}$ elliptic and $f\in L^{\infty}(0,1;L^{2}({\mathbb{R}}^{d}))$ such that the unique solution $u$ to (P) satisfies $u\notin L^{r}(0,1;L^{s}({\mathbb{R}}^{d}))$. Suppose that $u\in H^{\nu}(0,1;L^{2}({\mathbb{R}}^{d}))$. Since $u\in L^{2}(0,1;H^{1}({\mathbb{R}}^{d})$ by Theorem 2.1, complex interpolation (see [15, Theorem 14.7.12]) yields $u\in H^{\nu\theta}(0,1;H^{1-\theta}({\mathbb{R}}^{d}))$. By choice of $\theta$, the Sobolev embedding is applicable and gives $u\in L^{r}(0,1;L^{s}({\mathbb{R}}^{d}))$, a contradiction. ∎

### v2

@@ Abstract.
The first-named author is supported by the Alexander von Humboldt foundation by a Feodor Lynen grant. The second-named author is supported by NSF CAREER Grant DMS-2143668 and a Sloan Research Fellowship. The third-named author is supported by the VICI subsidy VI.C.212.027 of the Netherlands Organisation for Scientific Research (NWO)
@@ 1. Introduction
(CP) $\left\{\begin{aligned} u^{\prime}-\mathcal{A}u&=f,\ \ \text{on}\ (0,1),\\ u(0)&=0.\end{aligned}\right$.
@@
(P) $\left\{\begin{aligned} u^{\prime}-{\text{div}}(B\nabla u)&=f,\quad\text{on }(0,1),\\ u(0)&=0,\end{aligned}\right$.
@@
Acknowledgment
The authors thank the anonymous referee for their comments on the manuscript.
@@ 2.1. Known positive results
(2.1) $-\int_{0}^{1}\langle u(s),\phi^{\prime}(s)\rangle ds-\int_{0}^{1}\langle\mathcal{A}(s)u(s),\phi(s)\rangle ds=\int_{0}^{1}\langle f(s),\phi(s)\rangle ds.$
@@
2.2. A preliminary counterexample from the literature
[…]
(2.2) $\partial_{t}\zeta=\text{div}(B(t,\,x)\nabla\zeta)$
[…]
(2.3) $\zeta(t,\,x)=(1-t)^{-\mu/2}e^{-i\log(1-t)/2}w(x/(1-t)^{1/2}),$
(2.4) $B(t,\,x)=a(x/(1-t)^{1/2}),$
for appropriate choices of parameter $\mu\in\mathbb{R}$, function $w:\mathbb{R}^{d}\rightarrow\mathbb{C}$, and uniformly elliptic coefficients $a:\mathbb{R}^{d}\rightarrow\mathbb{C}^{d\times d}$. The parabolic PDE (2.2) is equivalent to the following elliptic PDE for $w$ on $\mathbb{R}^{d}$:
(2.5) $\text{div}(a(x)\nabla w)=\frac{1}{2}(iw+\mu w+x\cdot\nabla w).$
In [19] it is shown that for any choice of the parameter $\mu$ satisfying
(2.6) $0\leq\mu<d/2,$
one can find a uniformly elliptic matrix field $a$ on $\mathbb{R}^{d}$ that is smooth away from 0, and a solution $w$ to  (2.5) on $\mathbb{R}^{d}$ that is smooth away from 0 and locally Lipschitz on $\mathbb{R}^{d}$, such that the following estimates are satisfied: for all multi-indices $\alpha$ and all $|x|\geq 1$, we have
(2.7) $|\partial^{\alpha}w(x)|\leq C_{\alpha}|x|^{-|\alpha|-\mu},$
(2.8) $|\partial^{\alpha}a(x)|\leq C_{\alpha}|x|^{-|\alpha|}.$
The desired solutions $\zeta$ and coefficients $B$ for the parabolic problem (2.2) are then obtained through (2.3) and (2.4). We stress that the coefficients $a$ satisfy the boundedness and uniform ellipticity conditions
(2.9) $|a(x)\xi|\leq\Lambda|\xi|,\quad\text{Re}(a(x)\xi\cdot\overline{\xi})\geq\lambda|\xi|^{2}$
for some $\lambda,\,\Lambda>0$ depending on $\mu$ and all $x\in\mathbb{R}^{d}$ and $\xi\in\mathbb{C}^{d}$, so $B$ satisfies the conditions (1.1) with the same $\lambda$ and $\Lambda$.
We will only use the properties discussed above in the sequel. For the details of how $w$ and $a$ are chosen, the interested reader can consult Section 3 in [19]. Roughly speaking, for judicious choices of the form of $w$ and $a$, one can reduce the PDE (2.5) to a system of ODEs. These can be solved by fixing a choice of $w$ and solving for the coefficients in $a$.
Remark 2.3.
When split into real and imaginary parts, the equation (2.2) can be understood as a system of two equations, which are coupled through the imaginary part of $B$. An important philosophical point in [19] is that when the imaginary part of $B$ is taken to be symmetric, that part of $B$ doesn’t contribute to the ellipticity condition. This allows strong coupling of the equations without breaking the ellipticity condition.
@@ 2.3. Failure of $L^{r}(L^{s})$-integrability for variational solutions
The following counterexample is based on the construction from the previous subsection and is the basis for our subsequent counterexamples to Problems 1 and 2.
@@
Pick $\frac{2}{r}+\frac{d}{s}<\mu<\frac{d}{2}$ and fix corresponding $w$, $a$, $\zeta$, and $B$ as described in Section 2.2.
Step 1: general setup. It is clear from (2.2) and the properties of $w$ and $B$ that the weak derivative $\partial_{t}\zeta$ exists on $(0,T)$ as an $L^{2}(B_{R})$-valued function, for any $T<1$ and $R>0$, where $B_{R}$ is a shorthand notation for the Euclidean ball of radius $R$ in ${\mathbb{R}}^{d}$. Now let
@@
$\left\{\begin{aligned} \partial_{t}u(t,x)-{\text{div}}(B(t,x)\nabla u(t,x))&=f(t,x),\\ u(t,0)&=0.\end{aligned}\right$.
@@
By choice of $\mu$ we have $\frac{\mu+1}{2}+\frac{d}{4}>-\frac{1}{2}$, therefore also $\nabla u\in L^{2}(0,1;L^{2}({\mathbb{R}}^{d}))$.
@@
(2.11) $\begin{split}\|\partial_{m}\zeta(t,\cdot)\|_{L^{2}(B_{2}\setminus B_{1})}&=(1-t)^{-\frac{\mu}{2}-\frac{1}{2}}\|(\partial_{m}w)(\cdot/(1-t)^{1/2})\|_{L^{2}(B_{2}\setminus B_{1})}\\ &\leq C_{1}\||\cdot|^{-1-\mu}\|_{L^{2}(B_{2}\setminus B_{1})}.\end{split}$
@@
By choice of $\mu$, $r$ and $s$ we have $\frac{\mu}{2}+\frac{d}{2s}<-\frac{1}{r}$, therefore $u\not\in L^{r}(0,1;L^{s}({\mathbb{R}}^{d}))$. ∎
@@ 2.4. Negative result concerning problem 1
(2.12) $\left\{\begin{aligned} \partial_{t}u-{\text{div}}(B\nabla u)&=f,\\ u(0)&=0,\end{aligned}\right$.
@@
(2.13) $\left\{\begin{aligned} \partial_{t}\widetilde{v}-{\text{div}}(A\nabla\widetilde{v})&=\widetilde{g},\\ \widetilde{v}(-1)&=0,\end{aligned}\right$.
@@
$\int_{0}^{1}\langle g,u\rangle dt=\int_{-1}^{0}\langle\widetilde{g},\widetilde{u}\rangle dt$ $=\int_{-1}^{0}\langle(\widetilde{v})^{\prime},\widetilde{u}\rangle+((\widetilde{B})^{*}\nabla\widetilde{v}|\nabla\widetilde{u})_{2}dt$
@@
$\bigl|\int_{0}^{1}\langle g,u\rangle dt\bigr|$ $\leq\|v\|_{L^{p}(0,1;H^{1}({\mathbb{R}}^{d}))}\|f\|_{L^{p^{\prime}}(0,1;H^{-1}({\mathbb{R}}^{d}))}$
@@ 2.5. Negative result concerning problem 2
The strategy is as for the case $p>2$ in Theorem 2.5. Let $\nu\in(1/2,1]$ and let $0<\theta<\frac{1}{2\nu}$. Then define parameters $r$ and $s$ through $\frac{1}{r}=\frac{1}{2}-\nu\theta$ and $\frac{1}{s}=\frac{1}{2}-\frac{1-\theta}{d}$. They satisfy the condition in Theorem 2.5, consequently there is $B:[0,1]\times{\mathbb{R}}^{d}\to{\mathbb{C}}^{d\times d}$ elliptic and $f\in L^{\infty}(0,1;L^{2}({\mathbb{R}}^{d}))$ such that the unique solution $u$ to (P) satisfies $u\notin L^{r}(0,1;L^{s}({\mathbb{R}}^{d}))$. Suppose that $u\in H^{\nu}(0,1;L^{2}({\mathbb{R}}^{d}))$. Since $u\in L^{2}(0,1;H^{1}({\mathbb{R}}^{d}))$ by Theorem 2.1, complex interpolation (see [15, Theorem 14.7.12]) yields $u\in H^{\nu\theta}(0,1;H^{1-\theta}({\mathbb{R}}^{d}))$. By choice of $\theta$, the Sobolev embedding is applicable and gives $u\in L^{r}(0,1;L^{s}({\mathbb{R}}^{d}))$, a contradiction. ∎